THE MICROWAVE SPECTRUM OF HGeCl

\underline{WEI LIN}, Department of Chemistry, Wesleyan University, Middletown, CT 06459; LU KANG, Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138; STEW-ART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

HGeCl was produced by a high-voltage discharge of a free expansion of HGeCl_{3} diluted in argon. The rotational transitions were measured between 8 and 19 GHz by FTMW spectroscopy. The rotational constants and nuclear quadrupole coupling constants for various isotopomers were determined. The hyperfine constants will be compared with those of the related species, HSiCl.